Please help me with this: How much 12-karat (50%) gold must be mixed with pure gold to obtain 60 grams of 16-karat (66.6667%) gold? The mass of pure gold is not given.
Answer by mananth(16946)

How much 12-karat (50%) gold must be mixed with pure gold to obtain 60 grams of 16-karat (66.6667%) gold? The mass of pure gold is not given.
..
12-karat (50%) gold.............x grams
pure gold 100% ................60-x grams
to obtain 66.6667%............. 60 grams
...
50x+100(60-x)=66.6667*60
..
50x +6000-100x=4000 ( rounded off)
..
-50x=-6000+4000
-50x=-2000
/-50
x=40 grams of 50% gold is to be mixed.
...
m.annath@hotmail.ca